OpenStudy (chrisplusian):

someone please help!!!! "Find the radius of convergence of summation of 3(x-2)^n from n=0 to infinity

OpenStudy (chrisplusian):

this book comes directly from the textbook example. I am trying to understand how to do this. and here is my problem.......they use the ratio test and apply the limit as n approaches infinity. The catch is at the end they say "By the ratio test the series converges if the absolute value of (x-2) is less than one and diverges if the absolute value of (x-2) is greater than one. Therefore the radius of convergence is R=1. This statement doesnt make sense because if you plug one in for x in the absolute vqlue of x-2 then the statement says 1>1. I don't get it

OpenStudy (chrisplusian):

\[\lim_{n \rightarrow \infty}\left| \frac{ 3(x-2)^{n+1} }{ 3(x-2)^{n} } \right|=\lim_{n \rightarrow \infty}\left| x-2\right|\]

OpenStudy (chrisplusian):

this I understand

OpenStudy (anonymous):

the radius is half the length, but not the interval itself

OpenStudy (anonymous):

the 3 here is a red herring, it is just a number

OpenStudy (chrisplusian):

but then when you evaluate the limit you get \[\left| x-2 \right|\] if that is stictly less than one it converges by the ratio test. So: \[\left| x-2 \right| \lt 1\]

OpenStudy (chrisplusian):

so if you evaluate that you would see that x has to be stirctly less than three

OpenStudy (anonymous):

the limit is in \(n\) not in \(x\)

OpenStudy (chrisplusian):

how is it then that the radius is one?

OpenStudy (anonymous):

forget the \(x\), the limit is in \(n\)

OpenStudy (chrisplusian):

ok I am confused I thought that you could simplify and all the "n" s were gone.

OpenStudy (anonymous):

\(n\) is going to infinity, not \(x\) the possible values of \(x\) are the ones for which \(|x-2|<1\)

OpenStudy (anonymous):

yeah, the \(n\)'s are gone you are right

OpenStudy (anonymous):

so you have to make sure that that limit, which is \(|x-2|\) is less than one

OpenStudy (anonymous):

which is another way of saying that \(|x-2|<1\)

OpenStudy (chrisplusian):

I agree but when you evaluate that doesn't x have to be striclty less than three? not equal?

OpenStudy (anonymous):

if you solve \(|x-2|<1\) you get \(1<x<3\) an interval of length 2,

OpenStudy (chrisplusian):

agreed

OpenStudy (anonymous):

there is really no difference between \[\sum x^n\] and \[\sum(x-2)^n\] except you are shifted right two units

OpenStudy (chrisplusian):

but how is the radius then equal to one? wouldn't the radius of convergence have to be less than one?

OpenStudy (anonymous):

first one converges for \(|x|<1\) and the second one converges for \(|x-2|<1\)

OpenStudy (anonymous):

"radius" is half the length of the interval

OpenStudy (anonymous):

the word "radius' will make more sense in complex valued functions, because it is literally a "radius" but for real valued functions it is an interval on the real line

OpenStudy (chrisplusian):

I agree, I know it is a fine point, but it says that 1<x<3 which does not inlude 1 or 3, correct?

OpenStudy (anonymous):



OpenStudy (anonymous):

may include 1 or 3 or may not but that does not change the length of the interval

OpenStudy (anonymous):

in your case it does include 1 and does not include 3, as you can check but the inclusion or exclusion of the number at the endpoint of the interval does not change the length of the interval

OpenStudy (chrisplusian):

ok I understand the general idea, I just get hung up on the little points that seem trivial. Ok so we look for the interval and then check the endpoints?

OpenStudy (anonymous):

this is clear right? \((1,3)\) has the same length as \([1,3]\)

OpenStudy (chrisplusian):

Like two separate Ideas?

OpenStudy (anonymous):

just saying that the lengths are the same one point does not contribute anything to the length

OpenStudy (chrisplusian):

Oh yeah!!! that makes sense...... the interval notation made it click.... you should be a math professor, I would sign up for your calculus classes lol

OpenStudy (anonymous):

yeah they are separate ideas

OpenStudy (chrisplusian):

But then you probably would stop helping so much lol

OpenStudy (anonymous):

length of interval is one thing inclusion of the endpoints, one or the other or both, is separate

OpenStudy (chrisplusian):

thanks for your help.

OpenStudy (anonymous):

yw